Using Boyle's law and the value of the slope of the linearized relation P vs inverse V, show how one may calculate the number of molecules in the air column. How many molecules were in the air column of your apparatus?

The value of the slope was 2
Answer #1

The grapgh is P vs 1/V.

So, its slope m = P/(1/V) = PV (whose value is given as 2)

By Boyle's law, PV = constant = nRT

where n is the no. of moles, R is the gas constant and T is the temperature in Kelvin.

The eqn is PV = nRT,

we have the LHS of this eqn (=2), R = 8.314 J/(mol-K) (in SI unit so check if P and V are in SI unit too and if not conversion is needed) and n is required.

n = PV/RT = 2/(8.314*T)

T is the temperature at which the experiment is carried out. Plug in to get the numerical value of n.

n = no. of moles and 1 mole contains 6.022*10^23 molecules (Avogadro's number).